# Multiplexity amplifies geometry in networks

### v1

@@ Abstract
We analyzed 15 real-world multiplexes from various fields, whose details can be found in App. A. We investigated the overlap properties by pairing layers. That is, if we have a multiplex with $L=4$ layers, we study the overlap of $\binom{4}{2}=6$ unique pairs of layers. Then, we extracted the mutually connected component (MCC) of each pair of layers using the algorithm introduced in Ref. [26]. This ensures that, in each layer, there exists a path connecting the analyzed nodes, which is a necessary condition in our geometric model to define a meaningful notion of relative distance between them. It has also been shown that the MCC is relevant for many applications [27, 28, 29]. We will refer to the resulting mutually connected two-layer system as the pairwise MCC. Note that we only analyze these objects if they contain more than 100 nodes because smaller MCCs are prone to finite size effects. Once this procedure has been completed we construct the MG by keeping only those edges that are present in both layers of the pairwise MCC.
@@
First, the average degree $\langle\widetilde{k}\rangle$ of the MG, referred hereafter as the average mutual degree, measures the amount of edge overlap between two layers. We note that two sparse, independent Erdős-Renyi (ER) network realizations show next to no overlap. There, the probability of a link being present in both layers is $(\langle k\rangle/N)^{2}$, where $\langle k\rangle$ is the average degree and $N$ is the size of the network. This leads to a graph density in the MG that scales as $\sim N^{-2}$, or an average mutual degree of order $\sim N^{-1}$, resulting in very low values of both quantities in large scale networks. In fact, it was shown in Ref. [13] that these scalings hold for random graphs with arbitrary degree sequence. However, Fig. 1a, where we plot the average mutual degree as a function of the mean average degree of the two layers, shows that the amount of edge overlap in real multiplexes is much higher than the null-model would predict. This is in line with past results  [13, 16, 17] and with the intuition that there should be correlations between layers leading to similar connectivity patterns. We note, however, that the MG generally has an average degree much lower than the individual layers it is based on. In fact, $\langle\widetilde{k}\rangle<1$ for many pairwise MCCs.
@@
For a measure of the amount of overlapping triangles in the pairwise MCCs, we focus on the average local clustering coefficient $\langle\widetilde{c}\rangle$ of the corresponding MG, called hereafter the mutual clustering coefficient. In order for a triangle to be present in the MG, it needs to exist in both layers. Intuitively, this is highly unlikely in the case of independent graphs. In the Supplementary Information (SI) [30], we show that the mutual clustering coefficient scales as $\sim N^{-2}$ for sparse ER networks. However, Fig. 1b tells a different story. Here we plot the average local clustering of the MG against the mean of the two single-layer clustering coefficients. We see that not only does the mutual clustering not tend to zero, it remains of the same order as that of the single layer graphs. In some cases it even exceeds these levels. These results indicate that the correlations present between layers in multiplex networks have an especially strong effect on the triangle overlap.
@@
Figure 1: (a) The average degree of the MG versus the mean of the two average degrees of the single layer networks for several real multiplexes. One multiplex can result in several data points as multiplexes with more than two layers are deconstructed into all possible pairwise MCCs. The style of the data points is the same for all MCCs corresponding to the same multiplex, and we only label each type once. For example, MCCs related to the Malaria multiplex are represented by gray hexagons. The black dashed line indicates the diagonal $\langle\widetilde{k}\rangle=(\langle k^{(1)}\rangle+\langle k^{(2)}\rangle)/2$. (b) The analogous data for the average local mutual clustering coefficient. The details of the networks shown in this figure can be found in Appendix A
Clustering is an essential geometric graph property, linked to the triangle inequality inherent to all metric spaces. It is therefore appropriate to investigate the reported empirical observations through the lens of network geometry [23]. In particular, we adopt the similarity×$\times$×popularity picture [24, 31], in which nodes lie in an underlying similarity metric space modeled as a circle, where nearby vertices are thought to be more similar and therefore more likely to be connected. Simultaneously, some nodes are more popular than others, implying they will form more connections and end up with higher degrees [24].
@@
In the multiplex setting, this idea is formally implemented through the geometric multiplex model (GMM) [14] (see Appendix B for details). In the GMM, the different single layer networks making up the multiplex are embedded in an underlying similarity space, such that each node is assigned an angular coordinate $\theta^{(l)}_{i}$ in each layer and pairs of nodes are separated by similarity distances $d^{(l)}_{ij}=R\Delta\theta^{(l)}_{ij}$, where $\Delta\theta^{(l)}_{ij}=|\pi-|\pi-|\theta^{(l)}_{i}-\theta^{(l)}_{j}||$. The popularity dimension is incorporated by assigning to each node a hidden degree $\kappa^{(l)}_{i}$, which are chosen such that the observed degree distribution in a certain layer $l$ has a power law tail $P^{(l)}(k)\sim k^{-\gamma_{l}}$. The connection probability between two nodes is
@@
These steps imply that each layer can be seen as a realization of the single-layer geometric $\mathbb{S}^{1}$-model [24]. In the GMM, the correlations between the angular coordinates across layers are controlled through the parameter $g$. Perfectly correlated multiplexes will have equal coordinates in all layers and correspond to $g=1$. When $g=0$, the layers are uncorrelated. In this case, node’s coordinate in one layer are drawn independently from that node’s coordinates in other layers. Similarly, the correlations of hidden degrees are controlled through the parameters $\nu$. The parameter $\beta_{l}$ controls the geometric coupling strength of the $l$-th layer to its underlying similarity space, determining how strongly the distance between two nodes influences the probability of a connection between them. In the single-layer case, $\beta$ is the tuning parameter of a clustering phase transition [24]. In the strongly geometric regime, with $\beta>1$, clustering is finite. If $\beta\leq 1$, corresponding to the weakly geometric regime, it vanishes in the thermodynamic limit. Nevertheless, this decay is very slow, especially in the quasi-geometric regime $\beta_{c}^{\prime}<\beta\leq 1$ [25]. Here, the exact value of the transition point $\beta_{c}^{\prime}$ depends on the degree distribution.
@@
The first case is that of perfectly correlated layers ($g,\nu=1$) with homogeneous degree distributions. In the SI [30] we show analytically that, here, clustering scales with the system size as a power-law $\langle\widetilde{c}\rangle\sim N^{\sigma_{c}(\beta_{1},\beta_{2})}$. In Fig. 2a, we plot the dependence of the exponent $\sigma_{c}(\beta_{1},\beta_{2})\equiv\sigma_{c}$ on the geometric couplings of the two constituent networks. We can distinguish five regions based on whether the mutual clustering decays faster or slower than its single layer counterparts. In region I, $\beta_{1},\beta_{2}>1$ and so both layers are highly clustered. This behavior is carried over to the MG, where clustering remains finite, and so $\sigma_{c}(\beta_{1},\beta_{2})=0$. In regions II and III, one layer is either in the geometric or quasi-geometric regime, whereas the other is weakly geometric. The scaling behavior of the mutual clustering then lies in between that of the individual layers. In region IV, both layers are weakly geometric with $\beta_{1},\beta_{2}\leq 1$. However, the decay of the clustering coefficient is slower than that of the individual layers, implying the presence of significant levels of overlapping triangles. In particular, when $21\geq\beta_{1}=\beta_{2}\geq 1/$ 2, mutual clustering remains finite for all $N$ even though it vanishes in the large size limit of each separate layer. The fact that clustering decays very slowly in these regions is in line with the empirical results in Fig. 1b. Only when $\beta_{1}$ and $\beta_{2}$ are both very small (region V) does clustering decay fast. In fact, the scaling here is equivalent to that of the explicitly non-geometric case where $\beta_{1}=\beta_{2}=0$.
@@
Real networks are rarely homogeneous and we, therefore, investigate numerically the effect of degree heterogeneity on the scaling of the clustering coefficient, focusing on the region $\beta_{1}=\beta_{2}\equiv\beta$. In Fig. 2b, $\sigma_{c}$ is plotted against $\beta$ for $\gamma=2.5$, $\gamma=3.5$ and $\gamma=6$. We observe that the constant clustering for $\beta>1/2$ found in the homogeneous case is also present for heterogeneous networks. In fact, for $32<\gamma<$ 3 this region is extended all the way down to $\beta=0$. For networks with less pronounced heterogeneity ($\gamma>3$), the clustering does start to decrease for $\beta<1/2$. In the SI [30] we verify these results with analytic solutions for the explicitly non-geometric case $\beta=0$. The result of these calculations are shown in the inset of Fig. 2b. In Figs. 2c,d we explore the effect of imperfect correlations across layers ($g,\nu<1$). In all cases, reduced correlations lead to weaker mutual clustering.
The fact that both degree heterogeneity as well as geometry can lead to high clustering raises the question as to which effect is more important for the real multiplexes described in Fig. 1. In the SI [30] we randomize the initial layers while preserving their degree distribution, effectively decoupling them from their underlying metric spaces. We show that in almost all cases, this strongly decreases the mutual clustering, highlighting the importance of geometry. We also note that the decrease is less pronounced, or even absent, in the case of high degree heterogeneity, in line with our findings in Fig. 2b.
We perform a similar analysis for the average mutual degree $\langle\widetilde{k}\rangle$ in the SI [30]. Correlations in both the similarity as well as the popularity dimensions can lead to increased levels of edge overlap. However, the extended regions of constant overlap found for triangles are not observed in the case of edges. Only when both layers are strongly geometric ($\beta_{1,2}>1$) does one observe a constant average mutual degree. This is once again in line with our observations in real-world multiplexes, where edge-overlap is relatively low in comparison to that of triangles.
@@
To this end we employ the tool Mercator [33, 34], which uses a combination of machine learning and maximum likelihood techniques to faithfully embed networks into their hidden popularity×$\times$×similarity spaces. Here, we specifically leverage the tool’s estimation of the parameter $\beta$ from the observed clustering coefficient.
@@
Figure 3: The relation between the geometric couplings of the mutual graph and constituent single layer networks for several real multiplexes. White color corresponds to the transition point $\widetilde{\beta}=1$. In the inset, the effective coupling $\widetilde{\beta}$ is plotted against the smallest of the geometric couplings of the underlying layers.
@@
We first embedded several artificial mutual graphs constructed from pairs of $\mathbb{S}^{1}$ networks at varying $\beta_{1}$ and $\beta_{2}$ to obtain their effective geometric coupling $\widetilde{\beta}$. In order to obtain a fully connected MG, we set the average degrees of the individual layers to be relatively large ($\langle k^{(1)}\rangle=\langle k^{(2)}\rangle=50$). The results of this procedure are shown in Figs. 2e-h. There, the dotted line indicates the contour $\widetilde{\beta}=1$, which marks the transition point between the strongly and weakly geometric regimes. In panel (e) we focus on homogeneous degree distributions and perfect correlations. Notably, the supercritical regime $\widetilde{\beta}>1$ extends deep into the region where one or two of the constituent layers is weakly geometric. This occurs because the $\mathbb{S}^{1}$ connection probability is a decreasing function of the distance, meaning that nodes within individual layers predominantly form short-range connections. It is therefore much more likely that these are present in both layers. These short links are more closely aligned with the underlying metric space, which strengthens the effective geometric coupling, $\widetilde{\beta}$, and explains how weakly geometric layers can produce strongly geometric mutual graphs. This effect is strongest when the geometric couplings of the individual layers are equal. In App. B we show that, in this case, the geometric coupling in the GM is given by $\widetilde{\beta}\simeq\beta_{1}+\beta_{2}\equiv 2\beta$. The enhanced supercritical regime is robust to variations in the degree distribution, as evidenced in panel (f), which incorporates degree heterogeneity. Panels (g) and (h) show that while reducing correlations in the similarity dimension (e.g. setting $g=0.5$) leads to more weakly geometric mutual graphs, variations in the correlation strength within the popularity dimension have little effect.
@@
Finally, we performed the same analysis on the real-world multiplexes. We embedded the individual layers of their pairwise MCCs as well as the corresponding MG. Note that the MG for empirical networks generally has a low average degree, and might not be fully connected. In this case we only embed the giant connected component to obtain a value for $\widetilde{\beta}$. The behavior in Fig. 3 is in agreement with the model as the inferred effective coupling $\widetilde{\beta}$ is mostly high. In the inset of this figure, we see that generally $\widetilde{\beta}>1$, even when one of the two layers is weakly geometric. Of course, the increase in the geometricity of the empirical MGs is attenuated by the fact that real multiplexes do not have perfect correlations between the layers [14]. However, the enhanced geometry in these real multiplexes helps explain the high levels of triangle overlap and, to a lesser extent, the edge overlap reported in Fig. 1.
@@
In this paper, we have shown that edge and triangle overlap are higher than expected in many real multiplexes, with the effect being especially stark for triangles. We analyzed these results through the lens of network geometry, showing that the geometric multiplex model explains real-world observations. The edges that are most congruent with the underlying metric space are the ones present in all layers of the multiplex. This leads to an enhanced effective geometric coupling in the mutual graph, which is made up of the edges present in all layers. We confirm this effect in both synthetic and real-world multiplexes. Similar to multiplex percolation [35, 36] or the rich phase space of multilayer superconductor networks [37], multilayer-enhanced network geometry is a unique feature that is not present in the single-layer setting.

### v2

@@ Abstract
We analyzed 15 real-world multiplexes from various fields, whose details can be found in App. A. We investigated the overlap properties by pairing layers. That is, if we have a multiplex with $L=4$ layers, we study the overlap of $\binom{4}{2}=6$ unique pairs of layers. Then, we extracted the mutually connected component (MCC) of each pair of layers using the algorithm introduced in Ref. [26]. This ensures that, in each layer, there exists a path connecting the analyzed nodes, which is a necessary condition in our geometric model to define a meaningful notion of relative distance between them. It has also been shown that the MCC is relevant for many applications [27, 28, 29]. We will refer to the resulting mutually connected two-layer system as the pairwise MCC. Note that we only analyze these objects if they contain more than 100 nodes because smaller MCCs are prone to finite size effects. Once this procedure has been completed we construct the MG using the same set of nodes as in the pairwise MCC and by keeping only those edges that are present in both layers. Of course, we could define our mutual graph differently. For instance, one could keep edges as long as they were present in either of the two layers. However, our definition of the MG, which corresponds to the intersection of the two layers, is particularly useful for measuring overlap as defined in previous works [13, 16].
@@
First, the average degree $\langle\widetilde{k}\rangle$ of the MG, referred hereafter as the average mutual degree, measures the amount of edge overlap between two layers. We note that two sparse, independent Erdős-Renyi (ER) network realizations show next to no overlap. There, the probability of a link being present in both layers is $(\langle k\rangle/N)^{2}$, where $\langle k\rangle$ is the average degree and $N$ is the size of the network. This leads to a graph density in the MG that scales as $\sim N^{-2}$, or an average mutual degree of order $\sim N^{-1}$, resulting in very low values of both quantities in large scale networks. In fact, it was shown in Ref. [13] that these scalings hold for random graphs with arbitrary degree sequence. However, Fig. 1a, where we plot the average mutual degree as a function of the mean average degree of the two layers, shows that the amount of edge overlap in real multiplexes is much higher than the null-model would predict. This is in line with past results  [13, 16, 17] and with the intuition that there should be correlations between layers leading to similar connectivity patterns. We note, however, that the MG generally has an average mutual degree much lower than the individual layers it is based on. In fact, $\langle\widetilde{k}\rangle<1$ for many pairwise MCCs.
@@
For a measure of the amount of overlapping triangles in the pairwise MCCs, we focus on the average local clustering coefficient $\langle\widetilde{c}\rangle$ of the corresponding MG, called hereafter the mutual clustering coefficient. In order for a triangle to be present in the MG, it needs to exist in both layers. Intuitively, this is highly unlikely in the case of independent graphs. In the Supplementary Information (SI) [30], which includes Refs. [31, 25], we show that the mutual clustering coefficient scales as $\sim N^{-2}$ for sparse ER networks. However, Fig. 1b tells a different story. Here we plot the average local clustering of the MG against the mean of the two single-layer clustering coefficients. We see that not only does the mutual clustering not tend to zero, it remains of the same order as that of the single layer graphs. In some cases it even exceeds these levels. These results indicate that the correlations present between layers in multiplex networks have an especially strong effect on the triangle overlap.
These results points towards mutual graphs that are sparser than their corresponding individual layers while at the same time showing similar levels of clustering. In Fig. 1c we visualize this effect for two layers of the arXiv multiplex. We notice that the mutual graph indeed has a reduced amount of edges, even disintegrating into several disconnected components. At the same time, most nodes maintain or increase their local clustering coefficient with respect to the mean of the local clustering coefficient in the constituent graphs ($\Delta c_{i}=\widetilde{c}_{i}-(c^{(1)}_{i}+c^{2}_{i})/2\geq 0$).
@@
Figure 1: (a) The average mutual degree of the MG versus the mean of the two average degrees of the single layer networks for several real multiplexes. One multiplex can result in several data points as multiplexes with more than two layers are deconstructed into all possible pairwise MCCs. The style of the data points is the same for all MCCs corresponding to the same multiplex, and we only label each type once. For example, MCCs related to the Malaria multiplex are represented by gray hexagons. The black dashed line indicates the diagonal $\langle\widetilde{k}\rangle=(\langle k^{(1)}\rangle+\langle k^{(2)}\rangle)/2$. (b) The analogous data for the average local mutual clustering coefficient. The details of the networks shown in this figure can be found in Appendix A. (c) An arXiv pairwise MCC and its mutual graph. In the bottom two layers, representing the pairwise MCC, nodes are colored by the local clustering coefficient $c_{i}$. In the top layer, representing the MG, the nodes are colored by the difference between the local mutual clustering coefficient and the mean of the clustering of the corresponding nodes in the constituent layers $\Delta c_{i}=\widetilde{c}_{i}-(c^{(1)}_{i}+c^{(2)}_{i})/2$. In all layers, nodes with $k_{i}\leq 1$ are not shown as they do not contribute to the average local clustering coefficient.
[…]
@@
In the multiplex setting, this idea is formally implemented through the geometric multiplex model (GMM) [14] (see Appendix B for details). In the GMM, the different single layer networks making up the multiplex are embedded in an underlying similarity space, such that each node is assigned an angular coordinate $\theta^{(l)}_{i}$ in each layer and pairs of nodes are separated by similarity distances $d^{(l)}_{ij}=R\Delta\theta^{(l)}_{ij}$, where $|\Delta\theta^{(l)}_{ij}=|\pi-|\pi-|\theta^{(l)}_{i}-\theta^{(l)}_{j}$$|$. The popularity dimension is incorporated by assigning to each node a hidden degree $\kappa^{(l)}_{i}$, which are chosen such that the observed degree distribution in a certain layer $l$ has a power law tail $P^{(l)}(k)\sim k^{-\gamma_{l}}$. The connection probability between two nodes is
@@
These steps imply that each layer can be seen as a realization of the single-layer geometric $\mathbb{S}^{1}$-model [24]. In the GMM, the correlations between the angular coordinates across layers are controlled through the parameter $g$. Perfectly correlated multiplexes will have equal coordinates in all layers and correspond to $g=1$. When $g=0$, the layers are uncorrelated. In this case, a node’s coordinate in one layer is drawn independently from that node’s coordinats in another layer. Similarly, the correlations of hidden degrees are controlled through the parameters $\nu$. The parameter $\beta_{l}$ controls the geometric coupling strength of the $l$-th layer to its underlying similarity space, determining how strongly the distance between two nodes influences the probability of a connection between them. In the single-layer case, $\beta$ is the tuning parameter of a clustering phase transition [24]. In the strongly geometric regime, with $\beta>1$, clustering is finite. If $\beta\leq 1$, corresponding to the weakly geometric regime, it vanishes in the thermodynamic limit. Nevertheless, this decay is very slow, especially in the quasi-geometric regime $\beta_{c}^{\prime}<\beta\leq 1$ [25]. Here, the exact value of the transition point $\beta_{c}^{\prime}$ depends on the degree distribution.
@@
The first case is that of perfectly correlated layers ($g,\nu=1$) with homogeneous degree distributions. In the SI [30] we show analytically that, here, clustering scales with the system size as a power-law $\langle\widetilde{c}\rangle\sim N^{\sigma_{c}(\beta_{1},\beta_{2})}$. In Fig. 2a, we plot the dependence of the exponent $\sigma_{c}(\beta_{1},\beta_{2})\equiv\sigma_{c}$ on the geometric couplings of the two constituent networks. We can distinguish five regions based on whether the mutual clustering decays faster or slower than its single layer counterparts. In region I, $\beta_{1},\beta_{2}>1$ and so both layers are highly clustered. This behavior is carried over to the MG, where clustering remains finite, and so $\sigma_{c}(\beta_{1},\beta_{2})=0$. In regions II and III, one layer is either in the geometric or quasi-geometric regime, whereas the other is weakly geometric. The scaling behavior of the mutual clustering then lies in between that of the individual layers. In region IV, both layers are weakly geometric with $\beta_{1},\beta_{2}\leq 1$. However, the decay of the clustering coefficient is slower than that of the individual layers, implying the presence of significant levels of overlapping triangles. In particular, when $1\geq\beta_{1}=\beta_{2}\geq 1/2$, mutual clustering remains finite for all $N$ even though it vanishes in the large size limit of each separate layer. The fact that clustering decays very slowly in these regions is in line with the empirical results in Fig. 1b. Only when $\beta_{1}$ and $\beta_{2}$ are both very small (region V) does clustering decay fast. In fact, the scaling here is equivalent to that of the explicitly non-geometric case where $\beta_{1}=\beta_{2}=0$.
@@
Real networks are rarely homogeneous and we, therefore, investigate numerically the effect of degree heterogeneity on the scaling of the clustering coefficient, focusing on the region $\beta_{1}=\beta_{2}\equiv\beta$. In Fig. 2b, $\sigma_{c}$ is plotted against $\beta$ for $\gamma=2.5$, $\gamma=3.5$ and $\gamma=6$. We observe that the constant clustering for $\beta>1/2$ found in the homogeneous case is also present for heterogeneous networks. In fact, for $2<\gamma<3$ this region is extended all the way down to $\beta=0$. For networks with less pronounced heterogeneity ($\gamma>3$), the clustering does start to decrease for $\beta<1/2$. In the SI [30] we verify these results with analytic solutions for the explicitly non-geometric case $\beta=0$. The result of these calculations are shown in the inset of Fig. 2b. In Figs. 2c,d we explore the effect of imperfect correlations across layers ($g,\nu<1$). In all cases, reduced correlations lead to weaker mutual clustering.
The fact that both degree heterogeneity as well as geometry can lead to high clustering raises the question as to which effect is more important for the real multiplexes described in Fig. 1. In the SI [30] we perform degree preserving randomization [34] on the individual multiplex layers, effectively decoupling them from their underlying metric spaces. We show that in almost all cases, this strongly decreases the mutual clustering, highlighting the importance of geometry. We also note that the decrease is less pronounced, or even absent, in the case of high degree heterogeneity, in line with our findings in Fig. 2b.
We perform a similar analysis for the average mutual degree $\langle\widetilde{k}\rangle$ in the SI [30]. Correlations in both the similarity as well as the popularity dimensions can lead to increased levels of edge overlap. However, the extended regions of constant overlap found for triangles are not observed in the case of edges. Only when both layers are strongly geometric ($\beta_{1,2}>1$) does one observe a constant average mutual degree. This is once again in line with our observations in real-world multiplexes, where edge-overlap is relatively low in comparison to that of triangles. It also highlights the fact that high mutual clustering can exist even when edge overlap is low.
@@
[…]
@@
Figure 3: (a) The relation between the geometric couplings of the mutual graph and constituent single layer networks for several real multiplexes. White color corresponds to the transition point $\widetilde{\beta}=1$. (b) The effective coupling $\widetilde{\beta}$ is plotted against the smallest of the geometric couplings of the underlying layers. (c) The relationship between the average angular distance of connected nodes in the mutual graph and the mean of the average angular distances in the two corresponding single-layer MCCs.
@@
We first embedded several artificial mutual graphs constructed from pairs of $\mathbb{S}^{1}$ networks at varying $\beta_{1}$ and $\beta_{2}$ to obtain their effective geometric coupling $\widetilde{\beta}$. In order to obtain a fully connected MG, we set the average degrees of the individual layers to be relatively large ($\langle k^{(1)}\rangle=\langle k^{(2)}\rangle=50$). The results of this procedure are shown in Figs. 2e-h. There, the dotted line indicates the contour $\widetilde{\beta}=1$, which marks the transition point between the strongly and weakly geometric regimes. In panel (e) we focus on homogeneous degree distributions and perfect correlations. Notably, the supercritical regime $\widetilde{\beta}>1$ extends deep into the region where one or two of the constituent layers is weakly geometric. This occurs because the $\mathbb{S}^{1}$ connection probability is a decreasing function of the distance, meaning that nodes within individual layers predominantly form short-range connections. It is therefore much more likely that these are present in both layers. These short links are more closely aligned with the underlying metric space, which strengthens the effective geometric coupling, $\widetilde{\beta}$, and explains how weakly geometric layers can produce strongly geometric mutual graphs. This effect is strongest when the geometric couplings of the individual layers are equal. In the Supplementary Material [30] we show that, in this case, the geometric coupling in the GM is given by $\widetilde{\beta}\simeq\beta_{1}+\beta_{2}\equiv 2\beta$. The enhanced supercritical regime is robust to variations in the degree distribution, as evidenced in panel (f), which incorporates degree heterogeneity. Panels (g) and (h) show that while reducing correlations in the similarity dimension (e.g. setting $g=0.5$) leads to more weakly geometric mutual graphs, variations in the correlation strength within the popularity dimension have little effect.
@@
Finally, we performed the same analysis on the real-world multiplexes. We embedded the individual layers of their pairwise MCCs as well as the corresponding MG. Note that the MG for empirical networks generally has a low average mutual degree, and might not be fully connected. In this case we only embed the giant connected component (GCC) if this network has more than fifty nodes. The behavior in Fig. 3a is in agreement with the model as the inferred effective coupling $\widetilde{\beta}$ is mostly high. In Fig. 3b we see that generally $\widetilde{\beta}>1$, even when one of the two layers is weakly geometric. In Fig. 3c we show that it is mostly the short, geometric edges that exist in both layers of the multiplex, leading to a shorter average distance $\langle\Delta\theta\rangle_{E}$ between the connected nodes in the mutual graph than in the individual layers. Of course, the increase in the geometricity of the empirical MGs is attenuated by the fact that real multiplexes do not have perfect correlations between the layers [14, 37]. However, the enhanced geometry in these real multiplexes, where the mutual graph consists of mostly short, geometric edges, helps explain the high levels of triangle overlap reported in Fig. 1b.
@@
In this paper, we have shown that edge and triangle overlap are higher than expected in many real multiplexes, with the effect being especially stark for triangles. We analyzed these results through the lens of network geometry, showing that the geometric multiplex model explains real-world observations. The edges that are most congruent with the underlying metric space are the ones present in all layers of the multiplex. This leads to an enhanced effective geometric coupling in the mutual graph, which is made up of the edges present in all layers. We confirm this effect in both synthetic and real-world multiplexes.
We expect this enhanced geometry, and the resulting triangle overlap, to have significant implications for dynamical processes unfolding on such multiplex systems. Geometric correlations are known to influence a variety of dynamical processes, including navigability using protocols such as Greedy Routing (GR) [38] and the formation of Turing patterns in reaction-diffusion systems [39], as well as mitigating the extreme vulnerability of multiplex networks to targeted attacks [40]. Consequently, by controlling the geometric amplification and with it the levels of clustering in the multiplex, it is possible to influence transition thresholds, communicability, pattern formation, and network resilience. Furthermore, it was shown in Ref [20] that overlapping edges have a strong effect on spreading dynamics on multiplexes. These avenues will be explored in future work, within a broader context that investigates the interplay between dynamical processes and geometric networks. Finally, we note that, much like multiplex percolation [41, 42] or the rich phase space of multilayer superconductor networks [43], the emergence of multilayer-enhanced geometry is a genuinely new phenomenon—one that has no analogue in the single-layer setting.